OpenStudy (anonymous):

2a+4b=10 3a-b=1 Tell which equation you would use to isolate a variable. Explain the reasoning

OpenStudy (anonymous):

id recommend waiting for someone else's opinion, but, i would use 3a-b=1 if you were given one of the two variables to solve the equation, it would be simpler to divide once rather then multiple times to solve the equation for instance 3a-b=1 3a=b+1 a = (b+1)/3 the other formula 2a = 10 - 4b divide by 2 and a = 5 - 2b *Still new and relearning take this only as a method to check your work*

OpenStudy (anonymous):

I'd use the second one because it's more simple to isolate the b and you wouldn't have to divide by 10 and possibly end up with fractions. 3a-b=1 -b=-3a+1 ;minus 3a from both sides b= 3a-1 ;multiply by -1 to make b positive So b=3a-1 Now plug b into the first equation remember b=3a-1. So 2a+4(3a-1)=10 2a+12a-4=10 ;distribute 4 14a-4=10; combine like terms 14a=14; add 4 to both sides a=1; divide both sides by 14 So a=1

OpenStudy (anonymous):

thanks daydream you helped me as well, i knew you could "solve" these but its been years and i had forgotten the method

OpenStudy (anonymous):

yeah that happens with math, easy to forget. but yeah no problem

OpenStudy (anonymous):

thank you do much

OpenStudy (anonymous):

soo

OpenStudy (anonymous):

:) your welcome. keep trying, these become very simple